OpenStudy (anonymous):

the ratio of two numbers is 5/6 . the second number is 16 less than twice the first number. find both numbers. someone help! im stupid

OpenStudy (anonymous):

Let the first number be X. Second number is 16 less than twice: 2X - 16 = second number Ratio of two numbers = First number / second number = X / 2X - 16 = 5/6 Thus, 6X = 5(2X-16) 6X = 10X - 80 80 = 4X X = ? You're not stupid, you just haven't learned how to do it yet *NLP

OpenStudy (anonymous):

x=20 thank you